OpenStudy (anonymous):

Find the value of k that would make the function continuous in each case. g(x) = {(e^x-1)/x x=not0, k x=0}

OpenStudy (anonymous):

I'm sorry this is way too complicated @Yvonmei I wish I could help I'm so srry :(

OpenStudy (solomonzelman):

and, I don't see the k in any pieces of the function. Retype the question please.

OpenStudy (solomonzelman):

or is it kx=0, but then you didn't specify that it is for only one x-valuer, i.e. x=0

OpenStudy (solomonzelman):

I have typos, but most importantly fix the Q.

OpenStudy (kmeis002):

The function in question is this: \[ g_1(x) = \left \{ \begin{array}{lr} e^{x}-1: & x \neq 0 \\ k: & x=0 \end{array} \right.\] or: \[ g_2(x) = \left \{ \begin{array}{lr} e^{x-1}: & x \neq 0 \\ k: & x=0 \end{array} \right.\] Im going to assume it is the former (g1). In this case, for a function to be continuous, its limit must agree with the actual function value so: \[\lim_{x \to 0} f(x) = f(0) \implies \lim_{x \to 0} e^{x}-1 = k \implies 0=k\] So k must be zero

OpenStudy (kmeis002):

If that is the case then we get: \[\lim_{x \to 0} f(x) = f(0) \implies \lim_{x\to 0} \frac{e^x-1}{x} = k \] Note the limit shows an indeterminant form \( \frac{0}{0}\). Using L'hopital's rule we get: \[\lim_{x \to 0 } \frac{e^x -1}{x} = \lim_{x \to 0 } \frac{\frac{d}{dx} \left(e^x -1 \right ) }{\frac{d}{dx} x} = \lim_{x\to 0} = \frac{e^x}{1} = 1 = k\] So k = 1